OpenStudy (deadshot):

The function f(x)=sin2x contains which type of transformation or change to its parent graph?

OpenStudy (anonymous):

The parent graph is \(f(x)=\sin(x)\) So, in general, \(f(ax)=\sin(ax)\) will be horizontally compressed by a factor of \(\frac{1}{a}\)